OpenStudy (anonymous):

Vectors State if the lines are parallel or perpendicular or find the angle between the lines. x=1-3t y=2+4t z=-6+t and x=1+2s y=2-2s z=-6+s

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

<3,4,1> <2,-2,1> so they are noty parallel and not perpendicular also...

OpenStudy (anonymous):

What would make them parallel or perpendicular?

OpenStudy (anonymous):

cos theta=(3*2+4*(-2)+1)/(sqrt(26*9)=-0.065

OpenStudy (anonymous):

find the theta....

OpenStudy (anonymous):

if the direction cosines are same or multiplied of one then they are parallel...

OpenStudy (anonymous):

if lwo lines have direction cosines...<a ,b,c> and <d,e,f> then if a/d=b/e=c/f then the lines are parallel...if ad+be+cf =0 then parallel..

OpenStudy (anonymous):

Thanks for the help. Btw, the first vector has a minor typo and should be -3.

OpenStudy (anonymous):

yes.....sorry for that.

OpenStudy (anonymous):

np